OpenStudy (anonymous):

Which system of equations does this graph represent? Linear graph and parabola. They intersect at 2, negative 1 and negative 3, 4 y = x2 - 5 y = -x + 1 y = x2 - 5 y = -x - 1 y = x2 + 5 y = -x + 1 y = x2 + 5 y = -x - 1

OpenStudy (anonymous):

Would you like me to walk you through a solution, or are you just looking for an answer, because I won't just give answers and I don't like walking through these things when people don't stick around until the end and ask questions so. Let me know if you want real help.

OpenStudy (anonymous):

i want help

OpenStudy (anonymous):

If they intersect at (2, -1) and (-3, 4), these are the solutions to the system with a quadratic and linear equation in it.

OpenStudy (anonymous):



OpenStudy (animelover98):

What was the answer to this? Thank you! @cosmichaotic

OpenStudy (animelover98):

Ah nevermind, I figured it out. y = x2 - 5 y = -x - 1